If a fair coin is tossed 200 times what is the probability that you will see between 80 and 120 heads? If the Normal approximation to the Binomial distribution applies, use it. If not use the Binomial distribution methods.
Answer by stanbon(75887)

If a fair coin is tossed 200 times what is the probability that you will see between 80 and 120 heads? If the Normal approximation to the Binomial distribution applies, use it. If not use the Binomial distribution methods.
----------------
Use approx because np and nq are both greater than 10.
---
Find P(79.5 <= x <= 120.5)
---
mean = np = 200(1/2) = 100
std = sqrt(npq) = sqrt(100(1/2)) = 5sqrt(2)
-------------------
z(79.5) = (79.5-100)/5sqrt(2) = -5.798
---
z(120.5) = (120.5-100)/5sqrt(2) = 5.798
--------------
P(80 < x < 120) = P(-5.798 < z < 5.798) is approximately 1.0000
-------------
Using the binomial dist I get P(80< x <12) = 0.9964
=======================================================
Cheers,
Stan H.